OpenStudy (anonymous):

Andrew paid an agency 10% of his first year's rent for finding him an apartment. His rent is $1,800 per month. What is the amount of the commission that Andrew paid the agency?

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(1800\) is the rate of one month so what will be the rate of one year

OpenStudy (mathmath333):

so rate of one year =\(12\times 1800=21600\) what is \(10\%\) of \(21600\)

OpenStudy (anonymous):

To find 10% of the 21600 turn 10% into a decimal and times it b 21600 So, .10 x 21600 = ?

OpenStudy (anonymous):

it would be 216000

OpenStudy (mathmath333):

what is \(\dfrac{10}{100}\times 21600\)

OpenStudy (anonymous):

if you use a calculator it would be 21600 but you have to input your decimals

OpenStudy (mathmath333):

\(x\%=\dfrac{x}{100}\)

OpenStudy (anonymous):

2160?

OpenStudy (anonymous):

.10 is two places over so 216.00

OpenStudy (anonymous):

ohh true

OpenStudy (mathmath333):